# Radiative decays of $P$-wave bottom baryons from light-cone sum rules

### v1

@@ IV.4 The $[\mathbf{6}_{F},2,1,\lambda]$ doublet
Multiplets Baryon       Mass  Difference    Decay channels   Decay width     T.R.W.    T.S.W. Yang:2020zrh
($j^{P}$) (GeV) (MeV) (keV) (keV) (MeV)
  $[\mathbf{6}_{F},0,1,\lambda]$   $\Sigma_{b}({1\over 2}^{-})$ $6.05\pm 0.11$ $\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $62^{+150}_{-~{}62}$ $180^{+500}_{-180}$
$\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $120^{+350}_{-120}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $960^{+2600}_{-~{}960}$ $3500^{+9600}_{-3500}$ 1300
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $2500^{+7000}_{-2500}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $230^{+660}_{-230}$ $720^{+2100}_{-~{}720}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $490^{+1400}_{-~{}490}$
$[\mathbf{6}_{F},1,1,\lambda]$ $\Sigma_{b}({1\over 2}^{-})$ $6.06\pm 0.13$ $6\pm 3$ $\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $2.4^{+6.3}_{-2.4}$ $3.7^{+9.9}_{-3.7}$
$\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $1.3^{+3.6}_{-1.3}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $12^{+35}_{-11}$ $32^{+92}_{-31}$ $14.3^{+21.2}_{-10.9}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $20^{+57}_{-20}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $9.6^{+24.1}_{-~{}9.5}$ $14^{+38}_{-14}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $4.9^{+14.3}_{-~{}4.9}$
$\Sigma_{b}({3\over 2}^{-})$ $6.07\pm 0.07$ $\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $1.3^{+3.4}_{-1.3}$ $2.2^{+5.8}_{-2.2}$
$\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.86^{+2.35}_{-0.86}$
$\Sigma_{b}^{+}({3\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $21^{+53}_{-21}$ $34^{+91}_{-34}$ $4.8^{+5.9}_{-2.9}$
$\Sigma_{b}^{+}({3\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $13^{+38}_{-13}$
$\Sigma_{b}^{-}({3\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $5.4^{+13.3}_{-~{}5.4}$ $8.9^{+22.8}_{-~{}8.9}$
$\Sigma_{b}^{-}({3\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $3.5^{+9.5}_{-3.5}$
$[\mathbf{6}_{F},2,1,\lambda]$ $\Sigma_{b}({3\over 2}^{-})$ $6.11\pm 0.16$ $12\pm 5$ $\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $7.5^{+62.1}_{-~{}7.5}$ $11^{+92}_{-11}$ $51.4^{+76.5}_{-32.9}$
$\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $3.5^{+30.5}_{-3.5}$
$\Sigma_{b}^{+}({3\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $120^{+1100}_{-~{}120}$ $180^{+1600}_{-~{}180}$
$\Sigma_{b}^{+}({3\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $55^{+500}_{-~{}55}$
$\Sigma_{b}^{-}({3\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $31^{+260}_{-~{}31}$ $45^{+380}_{-~{}45}$
$\Sigma_{b}^{-}({3\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $14^{+120}_{-~{}14}$
$\Sigma_{b}({5\over 2}^{-})$ $6.12\pm 0.15$ $\Sigma_{b}^{0}({5\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $3.6^{+11.0}_{-~{}3.6}$ $3.6^{+11.0}_{-~{}3.6}$ $22.3^{+23.8}_{-13.8}$
$\Sigma_{b}^{+}({5\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $58^{+180}_{-~{}58}$ $58^{+180}_{-~{}58}$
$\Sigma_{b}^{-}({5\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $2.5^{+8.3}_{-2.5}$ $2.5^{+8.3}_{-2.5}$
$[\mathbf{6}_{F},1,0,\rho]$ $\Sigma_{b}({1\over 2}^{-})$ $6.05\pm 0.12$ $3\pm 1$ $\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $12^{+31}_{-12}$ $18^{+47}_{-17}$ $700$
$\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $6.2^{+17.6}_{-6.2}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $770^{+1900}_{-~{}760}$ $1200^{+3000}_{-1200}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $420^{+1100}_{-~{}420}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $200^{+500}_{-200}$ $300^{+770}_{-290}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $95^{+270}_{-~{}92}$
$\Sigma_{b}({3\over 2}^{-})$ $6.05\pm 0.12$ $\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $6.2^{+15.4}_{-~{}6.2}$ $14^{+37}_{-14}$ $410$
$\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $7.9^{+21.4}_{-~{}7.9}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $400^{+980}_{-400}$ $880^{+2300}_{-~{}880}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $480^{+1300}_{-~{}480}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $99^{+250}_{-~{}97}$ $220^{+560}_{-220}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $120^{+310}_{-120}$
@@
Multiplets Baryon       Mass    Difference    Decay channels   Decay width   T.R.W.   T.S.W. Yang:2020zrh
($j^{P}$) (GeV) (MeV) (keV) (keV) (MeV)
  $[\mathbf{6}_{F},0,1,\lambda]$   $\Xi_{b}^{\prime}({1\over 2}^{-})$ $6.20\pm 0.11$ $\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $180^{+410}_{-180}$ $580^{+1300}_{-~{}580}$ 1900
$\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $400^{+930}_{-400}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $420^{+1000}_{-~{}420}$ $740^{+1800}_{-~{}740}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $320^{+810}_{-320}$
  $[\mathbf{6}_{F},1,1,\lambda]$   $\Xi_{b}^{\prime}({1\over 2}^{-})$ $6.21\pm 0.11$ $7\pm 2$ $\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{{\prime}0}\gamma$ $10.0^{+8.3}_{-9.5}$ $15^{+19}_{-14}$ $4.7^{+5.8}_{-3.3}$
$\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $5.2^{+10.3}_{-~{}4.8}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{{\prime}-}\gamma$ $18^{+22}_{-18}$ $28^{+43}_{-28}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $10^{+21}_{-10}$
  $\Xi_{b}^{\prime}({3\over 2}^{-})$ $6.22\pm 0.11$ $\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{{\prime}0}\gamma$ $5.3^{+4.5}_{-5.2}$ $8.8^{+9.0}_{-8.7}$ $1.8^{+1.1}_{-1.0}$
$\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $3.5^{+4.5}_{-3.5}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{{\prime}-}\gamma$ $10.0^{+18.7}_{-~{}9.5}$ $17^{+31}_{-16}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $6.9^{+12.8}_{-6.4}$
  $[\mathbf{6}_{F},2,1,\lambda]$   $\Xi_{b}^{\prime}({3\over 2}^{-})$ $6.23\pm 0.15$ $11\pm 5$ $\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{{\prime}0}\gamma$ $19^{+44}_{-19}$ $27^{+73}_{-27}$ $27.3^{+28.5}_{-14.2}$
$\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $8.4^{+28.7}_{-~{}8.4}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{{\prime}-}\gamma$ $46^{+250}_{-~{}46}$ $67^{+340}_{-~{}67}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $21^{+86}_{-21}$
  $\Xi_{b}^{\prime}({5\over 2}^{-})$ $6.24\pm 0.14$ $\Xi_{b}^{\prime 0}({5\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $11^{+38}_{-11}$ $11^{+38}_{-11}$ $12.3^{+12.3}_{-~{}6.1}$
$\Xi_{b}^{\prime-}({5\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $26^{+92}_{-26}$ $26^{+92}_{-26}$
  $[\mathbf{6}_{F},1,0,\rho]$   $\Xi_{b}^{\prime}({1\over 2}^{-})$ $6.18\pm 0.12$ $3\pm 1$ $\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{{\prime}0}\gamma$ $32^{+72}_{-32}$ $48^{+120}_{-~{}48}$ $250$
$\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $16^{+45}_{-16}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{{\prime}-}\gamma$ $81^{+180}_{-~{}80}$ $140^{+330}_{-140}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $58^{+150}_{-~{}58}$
  $\Xi_{b}^{\prime}({3\over 2}^{-})$ $6.19\pm 0.11$ $\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $16^{+32}_{-15}$ $39^{+81}_{-36}$ $160$
$\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $23^{+49}_{-21}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{{\prime}-}\gamma$ $46^{+91}_{-44}$ $110^{+220}_{-100}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $60^{+130}_{-~{}56}$

### v3

@@ IV Decays of flavor $\mathbf{6}_{F}$ baryons
Multiplets Baryon       Mass  Difference    Decay channels   Coupling constants   Decay width     T.R.W.
($j^{P}$) (GeV) (MeV) ($\rm GeV^{-1}$) (keV) (keV)
  $[\mathbf{6}_{F},0,1,\lambda]$   $\Sigma_{b}({1\over 2}^{-})$ $6.05\pm 0.11$ $\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $0.22^{+0.12}_{-0.10}$ $62^{+150}_{-~{}62}$ $180^{+500}_{-180}$
$\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.25^{+0.16}_{-0.11}$ $120^{+350}_{-120}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $0.87^{+0.60}_{-0.36}$ $960^{+2600}_{-~{}960}$ $3500^{+9600}_{-3500}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $1.0^{+0.70}_{-0.40}$ $2500^{+7000}_{-2500}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $0.43^{+0.31}_{-0.17}$ $230^{+660}_{-230}$ $720^{+2100}_{-~{}720}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $0.50^{+0.35}_{-0.21}$ $490^{+1400}_{-~{}490}$
$[\mathbf{6}_{F},1,1,\lambda]$ $\Sigma_{b}({1\over 2}^{-})$ $6.06\pm 0.13$ $6\pm 3$ $\Sigma_{b}^{0}({1\over 2}^{-})\to\Lambda_{b}^{0}\gamma$ $2.20^{+0.64}_{-0.59}$ $36000^{+45000}_{-28000}$ $36000^{+45000}_{-28000}$
$\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $0.040^{+0.018}_{-0.013}$ $2.4^{+6.3}_{-2.4}$
$\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.024^{+0.010}_{-0.010}$ $1.3^{+3.6}_{-1.3}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $0.16^{+0.08}_{-0.05}$ $12^{+35}_{-11}$ $32^{+92}_{-31}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $0.094^{+0.040}_{-0.034}$ $20^{+57}_{-20}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $0.082^{+0.026}_{-0.022}$ $9.6^{+24.1}_{-~{}9.5}$ $14^{+38}_{-14}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $0.047^{+0.019}_{-0.016}$ $4.9^{+14.3}_{-~{}4.9}$
$\Sigma_{b}({3\over 2}^{-})$ $6.07\pm 0.07$ $\Sigma_{b}^{0}({3\over 2}^{-})\to\Lambda_{b}^{0}\gamma$ $2.90^{+1.00}_{-0.67}$ $63000^{+86000}_{-47000}$ $63000^{+86000}_{-47000}$
$\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $0.029^{+0.013}_{-0.011}$ $1.3^{+3.4}_{-1.3}$
$\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.035^{+0.015}_{-0.012}$ $0.86^{+2.35}_{-0.86}$
$\Sigma_{b}^{+}({3\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $0.12^{+0.05}_{-0.05}$ $21^{+53}_{-21}$ $34^{+91}_{-34}$
$\Sigma_{b}^{+}({3\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $0.14^{+0.05}_{-0.04}$ $13^{+38}_{-13}$
$\Sigma_{b}^{-}({3\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $0.058^{+0.026}_{-0.022}$ $5.4^{+13.3}_{-~{}5.4}$ $8.9^{+22.8}_{-~{}8.9}$
$\Sigma_{b}^{-}({3\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $0.070^{+0.031}_{-0.025}$ $3.5^{+9.5}_{-3.5}$
$[\mathbf{6}_{F},2,1,\lambda]$ $\Sigma_{b}({3\over 2}^{-})$ $6.11\pm 0.16$ $12\pm 5$ $\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $0.059^{+0.067}_{-0.037}$ $7.5^{+62.1}_{-~{}7.5}$ $11^{+92}_{-11}$
$\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.061^{+0.045}_{-0.030}$ $3.5^{+30.5}_{-3.5}$
$\Sigma_{b}^{+}({3\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $0.24^{+0.35}_{-0.15}$ $120^{+1100}_{-~{}120}$ $180^{+1600}_{-~{}180}$
$\Sigma_{b}^{+}({3\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $0.24^{+0.19}_{-0.10}$ $55^{+500}_{-~{}55}$
$\Sigma_{b}^{-}({3\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $0.12^{+0.14}_{-0.10}$ $31^{+260}_{-~{}31}$ $45^{+380}_{-~{}45}$
$\Sigma_{b}^{-}({3\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $0.15^{+0.10}_{-0.10}$ $14^{+120}_{-~{}14}$
$\Sigma_{b}({5\over 2}^{-})$ $6.12\pm 0.15$ $\Sigma_{b}^{0}({5\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.035^{+0.024}_{-0.022}$ $3.6^{+11.0}_{-~{}3.6}$ $3.6^{+11.0}_{-~{}3.6}$
$\Sigma_{b}^{+}({5\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $0.14^{+0.10}_{-0.10}$ $58^{+180}_{-~{}58}$ $58^{+180}_{-~{}58}$
$\Sigma_{b}^{-}({5\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $0.029^{+0.023}_{-0.012}$ $2.5^{+8.3}_{-2.5}$ $2.5^{+8.3}_{-2.5}$
$[\mathbf{6}_{F},1,0,\rho]$ $\Sigma_{b}({1\over 2}^{-})$ $6.05\pm 0.12$ $3\pm 1$ $\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $0.098^{+0.043}_{-0.028}$ $12^{+31}_{-12}$ $18^{+47}_{-17}$
$\Sigma_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.056^{+0.025}_{-0.015}$ $6.2^{+17.6}_{-6.2}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $0.39^{+0.17}_{-0.11}$ $770^{+1900}_{-~{}760}$ $1200^{+3000}_{-1200}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $0.23^{+0.10}_{-0.10}$ $420^{+1100}_{-~{}420}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $0.20^{+0.10}_{-0.10}$ $200^{+500}_{-200}$ $300^{+770}_{-290}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $0.11^{+0.05}_{-0.02}$ $95^{+270}_{-~{}92}$
$\Sigma_{b}({3\over 2}^{-})$ $6.05\pm 0.12$ $\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $0.070^{+0.028}_{-0.021}$ $6.2^{+15.4}_{-~{}6.2}$ $14^{+37}_{-14}$
$\Sigma_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.12^{+0.05}_{-0.04}$ $7.9^{+21.4}_{-~{}7.9}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{+}\gamma$ $0.28^{+0.11}_{-0.10}$ $400^{+980}_{-400}$ $880^{+2300}_{-~{}880}$
$\Sigma_{b}^{+}({1\over 2}^{-})\to\Sigma_{b}^{*+}\gamma$ $0.47^{+0.18}_{-0.12}$ $480^{+1300}_{-~{}480}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{-}\gamma$ $0.14^{+0.06}_{-0.04}$ $99^{+250}_{-~{}97}$ $220^{+560}_{-220}$
$\Sigma_{b}^{-}({1\over 2}^{-})\to\Sigma_{b}^{*-}\gamma$ $0.23^{+0.09}_{-0.06}$ $120^{+310}_{-120}$
@@
Multiplets Baryon       Mass    Difference    Decay channels   Coupling constants   Decay width   T.R.W.
($j^{P}$) (GeV) (MeV) ($\rm GeV^{-1}$) (keV) (keV)
  $[\mathbf{6}_{F},0,1,\lambda]$   $\Xi_{b}^{\prime}({1\over 2}^{-})$ $6.20\pm 0.11$ $\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $0.32^{+0.18}_{-0.12}$ $180^{+410}_{-180}$ $580^{+1300}_{-~{}580}$
$\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.37^{+0.21}_{-0.13}$ $400^{+930}_{-400}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.49^{+0.32}_{-0.20}$ $420^{+1000}_{-~{}420}$ $740^{+1800}_{-~{}740}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.33^{+0.22}_{-0.14}$ $320^{+810}_{-320}$
  $[\mathbf{6}_{F},1,1,\lambda]$   $\Xi_{b}^{\prime}({1\over 2}^{-})$ $6.21\pm 0.11$ $7\pm 2$ $\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{0}\gamma$ $2.60^{+0.61}_{-0.54}$ $44000^{+46000}_{-31000}$ $44000^{+46000}_{-31000}$
$\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $0.070^{+0.021}_{-0.020}$ $10.0^{+8.3}_{-9.5}$
$\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.040^{+0.011}_{-0.010}$ $5.2^{+10.3}_{-~{}4.8}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{-}\gamma$ $0.14^{+0.10}_{-0.10}$ $130^{+260}_{-110}$ $160^{+300}_{-140}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.096^{+0.040}_{-0.033}$ $18^{+22}_{-18}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.056^{+0.021}_{-0.021}$ $10^{+21}_{-10}$
  $\Xi_{b}^{\prime}({3\over 2}^{-})$ $6.22\pm 0.11$ $\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{0}\gamma$ $3.40^{+0.93}_{-0.74}$ $80000^{+88000}_{-55000}$ $80000^{+88000}_{-55000}$
$\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $0.050^{+0.016}_{-0.017}$ $5.3^{+4.5}_{-5.2}$
$\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.059^{+0.020}_{-0.016}$ $3.5^{+4.5}_{-3.5}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{-}\gamma$ $0.043^{+0.052}_{-0.051}$ $51^{+93}_{-51}$ $68^{+120}_{-~{}67}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.068^{+0.028}_{-0.024}$ $10.0^{+18.7}_{-~{}9.5}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.083^{+0.028}_{-0.028}$ $6.9^{+12.8}_{-6.4}$
  $[\mathbf{6}_{F},2,1,\lambda]$   $\Xi_{b}^{\prime}({3\over 2}^{-})$ $6.23\pm 0.15$ $11\pm 5$ $\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{{\prime}0}\gamma$ $0.10^{+0.12}_{-0.10}$ $19^{+44}_{-19}$ $27^{+73}_{-27}$
$\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.088^{+0.074}_{-0.037}$ $8.4^{+28.7}_{-~{}8.4}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{{\prime}-}\gamma$ $0.14^{+0.23}_{-0.10}$ $46^{+250}_{-~{}46}$ $67^{+340}_{-~{}67}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.14^{+0.14}_{-0.10}$ $21^{+86}_{-21}$
  $\Xi_{b}^{\prime}({5\over 2}^{-})$ $6.24\pm 0.14$ $\Xi_{b}^{\prime 0}({5\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.059^{+0.053}_{-0.035}$ $11^{+38}_{-11}$ $11^{+38}_{-11}$
$\Xi_{b}^{\prime-}({5\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.089^{+0.084}_{-0.054}$ $26^{+92}_{-26}$ $26^{+92}_{-26}$
  $[\mathbf{6}_{F},1,0,\rho]$   $\Xi_{b}^{\prime}({1\over 2}^{-})$ $6.18\pm 0.12$ $3\pm 1$ $\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{{\prime}0}\gamma$ $0.15^{+0.05}_{-0.07}$ $32^{+72}_{-32}$ $48^{+120}_{-~{}48}$
$\Xi_{b}^{\prime 0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.083^{+0.036}_{-0.028}$ $16^{+45}_{-16}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{{\prime}-}\gamma$ $0.24^{+0.10}_{-0.10}$ $81^{+180}_{-~{}80}$ $140^{+330}_{-140}$
$\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.16^{+0.06}_{-0.05}$ $58^{+150}_{-~{}58}$
  $\Xi_{b}^{\prime}({3\over 2}^{-})$ $6.19\pm 0.11$ $\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $0.10^{+0.04}_{-0.03}$ $16^{+32}_{-15}$ $39^{+81}_{-36}$
$\Xi_{b}^{\prime 0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.18^{+0.06}_{-0.05}$ $23^{+49}_{-21}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{{\prime}-}\gamma$ $0.17^{+0.06}_{-0.05}$ $46^{+91}_{-44}$ $110^{+220}_{-100}$
$\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.29^{+0.11}_{-0.10}$ $60^{+130}_{-~{}56}$